In how many ways can 4 prizes be distributed among 6 students, when(i) no student gets more than one prize?

Grade:11

1 Answers

aditya kulkarni
6 years ago
first of all we will select 4 students form the 6 and then we will distribute the prizes among the selected 4 students.
therefore total ways = 6C4(selection of 4 students) * 4!(number of arrangements)=15*4!=15*24=360
so there are total 360 ways